### v1

@@ 3.2.1 Decentralised Identity
A DID is a unique and persistent global identifier. DIDs enable verifiable and decentralized digital identities for entities that are under total control of the identity owner instead of relying on a central authority. A DID resolves to a DID document. The DID representation conforms with the Universal Resource Identifier (URI) scheme with three mandatory components as below
@@
$<did:robot:0xAb8483F64d9C6d1EcF9b849Ae677dD3315835cb2>$
The DID document possesses public keys for cryptographic operations, including authentication and verification. The DID document can also contain service endpoints, which can be used to advertise different services provided by the entity or the controller of the entity.
@@ 3.3 Verifiable Credentials (VC)
A VC represents statements made by an issuer in a tamper-evident and privacy-respecting manner. Specifically, VCs provide portable and provable claims about an entity that can be cryptographically verified. The exchange of VCs builds trust among DID-identified peers. The W3C maintains the VC specifications, which consist of three components: metadata, claims, and proof. Metadata describes meta-information about the VC, such as the issuer, expiration date, and public keys for verification. It also includes a set of claims about the entity and proofs.
@@ 3.4 Sovereign Robots
A sovereign robot is enabled with a Self Sovereign Identity (SSI). An identity that can be independently verified. Unlike centralised and federated identity management, SSI empowers robots with complete control over their identities, ensuring both absolute portability and enhanced security. The robot uses DID and VCs for communications and the establishment of claims.

### v2

@@ 3.2.1 Decentralised Identity
A Decentralized Identifier (DID) is a globally unique, cryptographically verifiable identifier that does not require a centralized registry or authority for its issuance or management. DIDs are a core component of decentralized identity systems and are defined by the W3C DID Specification. They are designed to enable self-sovereign identity—a model in which individuals or entities fully control their digital identifiers and associated data. A typical DID structure consists of the three mandatory components:-
@@
$<did:robo:0xAb8483F64d9C6d1EcF9b849Ae677dD3315835cb2>$
where, the meaning of the respective components is as follows:
  1. 1.
    $did$: The scheme indicating that it’s a decentralized identifier.
  2. 2.
    $method$ : Specifies the DID method, which defines how the identifier is created, resolved, and managed on a specific blockchain or distributed system.
  3. 3.
    $method-specifier$ : A unique string generated according to the method’s rules, often linked to a public key or hash.
In addition, a DID document serves as a metadata container that describes the cryptographic and service-related capabilities associated with a specific decentralized identity (DID). It typically includes one or more public keys used for cryptographic operations such as authentication, authorisation, and data integrity verification. These keys enable secure interactions between entities without relying on centralized intermediaries. In addition to public keys, the DID document may define one or more service endpoints, URLs, or network addresses that facilitate interactions with services controlled by the DID subject or its controller. These endpoints can support use cases such as credential issuance, verification services, messaging, or decentralized data exchange. The structure and semantics of the DID document conform to the W3C DID specification, ensuring interoperability across different platforms and implementations.
@@ 3.3 Verifiable Credentials (VC)
A verifiable credential (VC) is a tamper-evident, cryptographically secure digital assertion made by an issuer about a subject. It includes three key components: ”metadata,” which provides contextual details (e.g., issuer, issuance date, and public key references); ”claims,” which are statements about the subject (e.g., qualifications or certifications); and ”proof,” a cryptographic signature ensuring authenticity and integrity. VCs enable trustless, portable, and privacy-preserving information exchange in decentralized identity systems, supporting ”self-sovereign identity” (SSI). They allow secure, offline verification without relying on centralized authorities, while also enabling features like selective disclosure and zero-knowledge proofs for enhanced privacy.
@@ 3.4 Sovereign Robots
A sovereign robot is equipped with a Self-Sovereign Identity (SSI)—a decentralised, verifiable digital identity that operates independently of centralized or federated identity providers. SSI grants the robot full autonomy over the creation, control, and management of its identity, ensuring maximum portability, resilience, and security. Leveraging Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs), the robot can establish cryptographically verifiable claims and engage in trusted, peer-to-peer communications with other entities in a decentralized ecosystem.